INAF queue observing, 2020 Jan 29 - Jan 30

Observers: A. Rossi, S. Paiano & Telescope Operator: S. Allanson & Support Astronomer: D. Thompson & Instrument: MODS1-2

At sunset the sky is partly cloudy but it is clearing up

Summary

Night duration (12 degree - twilight) 1:30 - 13:30 UT = 12h 00m = 43200 seconds
Observing time: 10h 50min (90%)

Weather time loss 45min (6%) sec

Technical time loss: 25min (4%) sec

Data Summary

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images NotesSorted ascending
Antoniucci IRAS06+10 MODS1-2 dual, slit 0.8" 0.6/0.3h mods1r.0053-56,mods1b.0023-26 mods2r.0052-5,mods2b.0024-27 completed
Mannucci J0415+16 MODS1-2 dual, 0.6"   mods1r.0026,36, ,mods1b.0019,22 mods2r.0026,36 ,mods2b.0020,23 completed
RossiDDT SN2018ibb MODS1-2 dual, 1.2" 1h/3h mods1r.0030-31,mods1b.0020-21 mods2r.0030-31,mods2b.0021-22 completed
Belladitta PS0J0944 MODS1-2 red, 1.2" 2.8/2.6h mods1r.0040-0046, mods2r.0040-0045 completed
Calibrations slit flats MODS1-2 dual, slit 1"     for Wolter
Calibrations flats MODS1-2 dual grating, z     imaging z flats
Schmidt J1149+2813 LBC red r i z /1.75,0.75,0.25   LBCr only
Calibrations slit flats MODS1-2 red, slit 5"     STD Belladitta
standard G191 MODS1-2 dual, slit 5",   mods1r.0019-0020,mods1b.0015-16 mods2r.0019-20,mods2b.0016-17 STD for Paiano
standard G191 MODS1-2 dual, slit 1.2"   mods1r.0021-0022,mods1b.0017-18 mods2r.0021-22,mods2b.0018-19 STD for Paiano

Details

MODS CALIBRATIONS for tomorrow missing:

01:42 Slew to STD G191 dual grating
  • acq mods1r.20200130.0016-0018, mods2r.20200130.0016-0018
  • sci mods1r.0019-0020,mods1b.0015-16 mods2r.0019-20,mods2b.0016-17 dual 5
  • sci mods1r.0021-0022,mods1b.0017-18 mods2r.0021-22,mods2b.0018-19 dual 1.2"

02:18 Slew to Mannucci program, target J0415
  • acq mods1r.0023-0025, mods2r.0023-0025
  • DIMM seeing 1.5" variable
  • sci mods1r.0026,mods1b.0019 mods2r.0026,mods2b.0020
  • seeing down to 0.9 at the end
  • we noticed that in MODS1 the slit is slighty tilted, i.e. not exactly vertical respect to the detector

02:43 Slew to Rossi DDT program, target SN2018ibb
  • acq mods1r.0027-0029, mods2r.0027-0029
  • DIMM seeing 0.9"
  • sci mods1r.0030-31,mods1b.0020-21 mods2r.0030-31,mods2b.0021-22

03:33 Slew to Mannucci program, target J0415
  • acq mods1r.0032-0035, mods2r.0032-0035
  • DIMM seeing 1.0"
  • sci mods1r.0036,mods1b.0022 mods2r.0036,mods2b.0023-

03:58 Slew to Belladitta, target PS0J0944 +83
  • DIMM seeing 2.0" with some passing cirrus
  • acq mods1r.0037-0039, mods2r.0037-0039
  • sci mods1r.0040-0046, mods2r.0040-0045
  • 4th exposure on mods2r is lost for unknown reason (mods2 was frozen, after red expdone it jumped to exp #5)
  • 13 min loss
  • we take another 780 exp
  • seeing 0.9", but huumidity 90%

06:00 Slew to Mannucci program,

06:05 close due to humidity 95%

06:45 The humidity dropped and we can open again the telescope. Unfortunately we have some problems with the start-up of LBC. Investigations are on-going.

07:30 There are newtork and communication problem with the systems of LBC because the power supply is dead. There is not the possibility to recover it during the night. Finally we lost ~45 minutes for humidity.

07:35 For this reason, we decide to switch again to MODS.

08:03 We are moving on the Antoniucci program and we are pointing IRAS06297+10
  • DIMM seeing 1.1"
  • acq mods1r.0050-0051, mods2r.0049-0050
  • sci mods1r.0053-56,mods1b.0023-26 mods2r.0052-5,mods2b.0024-27

08:45 We are changing instrument. We use only LBC-Red. LBC-Blue has communication problem. Can be solved only during day time
  • 10 min loss

09:30 Slew to Schimidt field J1149+2813,
  • Red Only OB created with r i z imaging
  • seeing 1" on image
  • dofpia, X2, RedOnly
  • copointing with ptautoadjust
  • 09:40 redonly_riz_1
  • 0.8" on image at 09:50
  • dofpia, RedOnly
  • 10:15 redonly_riz_2
  • we are very high, so due to high rotation collimation is not perfect: so the next 20 min data have star not perfectly round
  • 1.2-1.5" on image at 10:35
  • dofpia cannot be run, we are to high, but was not so bad (because seeing is not good) let's tae few more images
  • redonly_riz_2 (repeat the OB#2)
  • 1.0-1.15" on image at 11:12
  • dofpia, RedOnly
  • 11:23 redonly_riz_3
  • 0.9-0.95" on image at 11:34
  • 0.9-1.1" on image at 11:42
  • dofpia, RedOnly
  • 11:55 redonly_riz_4
  • 0-7"-0.9" on image at 12:05
  • ~1.05" on image at 12:18 and 12:25
  • dofpia, RedOnly at 12:25
  • 0.7" on image at 12:48
  • 12:36 redonly_riz_6
  • 13:06 redonly_riz_7
  • redonly_riz_5 and 8 TBD

CALIBRATIONS
  • SKYFLATS LBC Uspec e r-sloan
  • bias

-- AndreaRossi - 30 Jan 2020
Topic revision: r1 - 31 Jul 2024, UnknownUser
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback